# Optimal performance of a three stroke heat engine in the microscopic regime

### v1

@@ III.2.1 Optimal performance of the engine for finite sized heat baths
This suggests that when a two-level system functions as a heat bath, it restricts thermal operations between an identity map and thermalization. As a result, when we model a two-level system resonant with the working body as a heat bath, the three stroke engine cannot extract work since the heat stroke fails to induce non-passivity in the working body.
@@
Similarly, in this scenario, one can obtain an expression for optimal work production ($W_{d}$) by substituting the values of $\lambda$s given in Eq. (45) into Eq. (29). We plot the work production and efficiency in Figs. 2 and 3 for $\omega=0.2/\beta_{H}$ and some $d$ values, and compared them with $W_{\max}$ and $\eta_{\max}$, respectively.
@@ III.2.2 Optimal performance of the engine modelled with Jaynes-Cummings interaction
In this section, we explore the Jaynes-Cummings model and determine the optimal work and efficiency for the stroke-based discrete engine we considered. Using Eq. (22), we obtain the maximum achievable values of $\lambda$s for hot and cold heat baths as follows[65]:
@@
Figure 4: Plot of efficiencies vs $\beta_{C}/\beta_{H}$ with $\omega=0.2/\beta_{H}$. The solid, dashed, dotdashed, and dotted curves represent Carnot efficiency, optimal efficiency, efficiency for finite heat bath with $d=5$, and efficiency for Jaynes-Cummings model, respectively.
Figure 5: Plot of $\beta_{H}$ork vs Efficiency. The solid, dotdashed, dashed, and dotted curves represent optimal work, work for finite heat bath with $d=10$, $d=5$, and work for Jaynes-Cummings, respectively.
Now, substituting $\lambda^{\textrm{max}}_{H,\text{JC}}$ and $\lambda^{\textrm{max}}_{C,\text{JC}}$ from Eqs. (47) and (48) into Eqs. (29) and (30), we can obtain the optimal values of work and efficiency for the engine. The resulting expressions are cumbersome, hence we opted not to put them here; rather, we compare the efficiency obtained in this case with the other cases as shown in Fig. 4. Furthermore, we also make a comparison of the tradeoff between work and efficiency in Fig. 5 for various scenarios we have discussed so far.
@@ IV Conclusion
Investigating the significance of coherence in the operation of three-stroke engines presents a captivating avenue for research. In Ref. [66], it was shown that coherence in the state of the working body, modeled as a two-level system, fails to yield performance enhancements in three-stroke engines. This raises the question: is it possible to generalize this result? Specifically, it will be interesting to prove or disprove that the coherence present in the state of the working body is not helpful to enhance the performance of a three-stroke engine. However, characterizing the transformation of coherence in a state under thermal operations remains a formidable challenge, particularly beyond two dimensions. Nonetheless, related results suggest that the dephasing of the state in the local energy eigenbasis is compatible with thermal operations across all dimensions. This result might be helpful in this endeavor.
@@ Appendix A Upper bounding the efficiency of the three stroke engine using Carnot efficiency
$Q_{C}=\text{Tr}\Big{(}H_{C}\big{(}{\cal E}_{C}(V{\cal E}_{H}(\rho_{S})V^{\dagger})-V{\cal E}_{H}(\rho_{S})V^{\dagger}\big{)}\Big{)}$
$\,\,\,\,\,\,\,\,\,\,=\text{Tr}\Big{(}H_{C}\big{(}\rho_{S}-V{\cal E}_{H}(\rho_{S})V^{\dagger}\big{)}\Big{)}$ (53)
@@ Appendix B Thermal operation and $\beta$-ordering
where $U$ is a joint unitary acting on the system and a heat bath that commutes with the total Hamiltonian of the system and bath, i.e., $[U,H_{S}+H_{E}]=0$, and $\gamma_{E}$ is a thermal state of the bath with very large volume at some fixed inverse temperature $\beta$. If $\rho_{S}$ is an energy-incoherent state, i.e., $[\rho_{S},H_{S}]=0$, one can prove [26]
@@ Appendix C Proof of theorem 2
Now, theorem 6 of Ref. [20] states that the optimal work output and efficiency for an open cycle engine are obtained at the extremal points of the thermal cone $\mathcal{T}_{H}(\bm{p})$, where the work extracting unitary in the work stroke is provided by
@@
during the heat stroke. After the work stroke which is governed by the unitary $X$ in Eq. (72), the resulting population vector from the heat stroke given in Eq. (77) transformed to
@@ Appendix D Range of thermal operation with finite sized baths
Note that the energies of the blocks from left to right are $0,\omega,2\omega,\ldots,(d+1)\omega$. As the unitary and the initial state of the system and bath have similar block diagonal structures, $U\left(\rho_{S}\otimes\rho_{E}\right)U^{\dagger}$ also has a similar block diagonal structure. Therefore, it is enough to find out the elements of one block. Consider a block position $j$ with energy $j\omega$, then we have

### v2

@@ III.2.1 Optimal performance of the engine for finite sized heat baths
This suggests that when a two-level system functions as a heat bath, it restricts thermal operations between an identity map and thermalization. As a result, when we model a two-level system resonant with the working body as a heat bath, the three stroke engine cannot extract work since the heat stroke fails to induce non-passivity in the working body. Observe that Eq. (43) resembles Eq. (50). This suggests that the process of re-thermalizing the state of the two-level working body for an open-cycle engine (discussed in Sec. II.3) is identical to the proposed three-stroke engine that interacts with a cold heat bath modeled using a truncated harmonic oscillator with two levels (this can be observed by substituting $d=1$ in Eq. (48)).
@@
Similarly, in this scenario, one can obtain an expression for optimal work production ($W_{d}$) by substituting the values of $\lambda$s given in Eq. (52) into Eq. (LABEL:Opt_performance_W). We plot the work production and efficiency in Figs. 2 and 3 for $\omega=0.2/\beta_{H}$ and some $d$ values, and compared them with $W_{\max}$ and $\eta_{\max}$, respectively.
@@ III.2.2 Optimal performance of the engine modelled with Jaynes-Cummings interaction
In this section, we explore the Jaynes-Cummings model and determine the optimal work and efficiency for the stroke-based discrete engine we considered. Let us recall that the Hamiltonian of the system is $H_{S}=\omega|1\rangle\!\langle 1|$. The Hamiltonian of the bath is modelled by a single mode harmonic oscillator, i.e.,
$H_{E}=\sum_{n=0}^{\infty}n\omega|n\rangle\!\langle n|,$ (54)
where $|{n}\rangle$ denotes the eigenstate of $H_{E}$ with eigenvalue $n\omega$. The interaction between the working body and the bath is described by a resonant Jaynes-Cummings Hamiltonian $H_{\text{JC}}$ in rotating wave approximation as follows:
$H_{\text{JC}}=g(\sigma_{+}\otimes a+\sigma_{-}\otimes a^{\dagger}),$ (55)
where $a^{\dagger}$ and $a$ represent the creation and annihilation operators on the bath, while $\sigma_{+}=|1\rangle\!\langle 0|$ and $\sigma_{-}=|0\rangle\!\langle 1|$ represent the excitation and de-excitation operators on the working body, respectively. Clearly $[H_{S}+H_{E},H_{\text{JC}}]=0$, therefore the transformation
$\rho_{S}$$\rightarrow$ $\text{Tr}_{E}\left(e^{-itH_{\text{JC}}}\left(\rho_{S}\otimes\frac{e^{-\beta_{H}H_{E}}}{\text{Tr}\left(e^{-\beta_{H}H_{E}}\right)}\right)e^{itH_{\text{JC}}}\right)$ (56)
::=:= ${\cal E}_{H,\text{JC}}\left(\rho_{S}\right),$
is a thermal operation at the inverse temperature $\beta_{H}$, which we denoted by $\mathcal{E}_{H,\text{JC}}$. In a similar manner, starting from the initial state $V\mathcal{E}_{H,\text{JC}}(\rho_{S})V^{\dagger}$, one can describe the thermal operation $\mathcal{E}_{C,\text{JC}}$ in the third stroke at the inverse temperature $\beta_{C}$ (recall that from Eq. (15), thermal operation ${\cal E}_{C,\text{JC}}$ in the third stroke acts on the state that results after the second stroke, i.e., $V\mathcal{E}_{H,\text{JC}}(\rho_{S})V^{\dagger}$). The maximum achievable values of $\lambda^{\text{max}}_{H}$ and $\lambda^{\text{max}}_{C}$ (denoted by $\lambda^{\text{max}}_{H,\text{JC}}$ and $\lambda^{\text{max}}_{C,\text{JC}}$) for the thermal operations ${\cal E}_{H,\text{JC}}$ and ${\cal E}_{C,\text{JC}}$ in the first and third strokes when it is modelled by Jaynes Cummings interaction are as follows: (see Eq. 10 of Ref. [63])
@@
Figure 4: Plot of efficiency ($\eta$) vs $\beta_{C}/\beta_{H}$ with $\omega=0.2/\beta_{H}$. The dashed, solid, dotdashed, and big-dashed curves represent Carnot efficiency, optimal efficiency, efficiency for finite heat bath with $d=10$, and efficiency for Jaynes-Cummings model, respectively.
Figure 5: Plot of $\beta_{H}W$ vs efficiency ($\eta$), where $W$ represents work done by the engine. The solid, dotdashed, dotted, and big-dashed curves represent optimal work, work for finite heat bath with $d=10$, $d=5$, and work for Jaynes-Cummings, respectively.
Now, substituting $\lambda^{\textrm{max}}_{H,\text{JC}}$ and $\lambda^{\textrm{max}}_{C,\text{JC}}$ from Eqs. (57) and (58) into Eqs. (LABEL:Opt_performance_W) and (34), we can obtain the optimal values of work and efficiency for the engine. The resulting expressions are cumbersome, hence we opted not to put them here; rather, we compare the efficiency obtained in this case with the other cases as shown in Fig. 4. Furthermore, we also make a comparison of the tradeoff between work and efficiency in Fig. 5 for various scenarios we have discussed so far.
@@ IV Conclusion
Investigating the significance of coherence in the operation of three-stroke engines presents a captivating avenue for research. In this paper, we have focused only on the working bodies that do not have any coherence in energy eigenbasis (i.e., in the eigenasis of Hamiltonian $H_{S}$). Ref. [52] demonstrated that coherence in the state of the working body, modeled as a two-level system, fails to improve the performance of a three-stroke engine. This raises the question: can this result be generalized for arbitrary dimensional working body? Specifically, it will be interesting to prove or disprove that the coherence present in the state of the working body is not useful to improve the performance of a three-stroke engine. However, characterizing the transformation of coherence in a state under thermal operations remains a formidable challenge, particularly beyond two dimensions. Nonetheless, related results suggest that the dephasing of the state in the local energy eigenbasis is compatible with thermal operations in all dimensions [56, 54]. These results might be helpful in this endeavor.
@@ Appendix A Upper bounding the efficiency of the three stroke engine using Carnot efficiency
$Q_{C}=\text{Tr}\Big{(}H_{S}\big{(}{\cal E}_{C}(V{\cal E}_{H}(\rho_{S})V^{\dagger})-V{\cal E}_{H}(\rho_{S})V^{\dagger}\big{)}\Big{)}$
$\,\,\,\,\,\,\,\,\,\,=\text{Tr}\Big{(}H_{S}\big{(}\rho_{S}-V{\cal E}_{H}(\rho_{S})V^{\dagger}\big{)}\Big{)}$ (63)
@@ Appendix B Thermal operation and $\beta$-ordering
where $U$ is a joint unitary operation acting on the system and a heat bath that commutes with the total Hamiltonian of the system and bath, i.e., $[U,H_{S}+H_{E}]=0$, and $\gamma_{E}$ is a thermal state of the bath with very large volume at some fixed inverse temperature $\beta$. If $\rho_{S}$ is an energy-incoherent state, i.e., $[\rho_{S},H_{S}]=0$, one can prove [27]
@@ Appendix C Proof of theorem 2
Now, theorem 6 of Ref. [20] states that the optimal work output and efficiency for an open cycle engine are obtained at the extremal points of the thermal cone $\mathcal{T}_{H}(\bm{p})$, where the unitary operation in the work stroke is provided by
@@
during the heat stroke. After the work stroke which is governed by the unitary operation $X$ in Eq. (82), the resulting population vector from the heat stroke given in Eq. (87) transformed to
@@ Appendix D Range of thermal operation with finite sized baths
Note that the energies of the blocks from left to right are $0,\omega,2\omega,\ldots,(d+1)\omega$. As the unitary operation $\left(U\right)$ and the initial state of the system and bath have similar block diagonal structures, $U\left(\rho_{S}\otimes\rho_{E}\right)U^{\dagger}$ also has a similar block diagonal structure. Therefore, it is enough to find out the elements of one block. Consider a block position $j$ with energy $j\omega$, then we have